OpenStudy (blackops2_143):

Which answer describes the solution to each equation? Column A: 1. 2x + 6 - x = 1/3 (3x +18) 2. 1/8x + 7 = 19 3. 3x+3 OVER 3 = X + 1 4. 3+7x + 6 = 2x + 5 + 5x Column B: A. one solution B. no solution C. an infinite number of solutions Match A,B,and C to the equations

OpenStudy (calculusxy):

Do you know what one solution, no solution, and an infinite number of solutions mean?

OpenStudy (blackops2_143):

yes

OpenStudy (welshfella):

something like x - 2 = x - 2 will have infinite solutions you can plug in infinte values of x into it and they will satissfy the thing

OpenStudy (calculusxy):

Then solve for the variable and check whether it has no solution, infinite number, or only one.

OpenStudy (calculusxy):

For example, I will do number one and check whether this fits either one of the three answers. \(2x + 6 -x = \frac{1}{3}(3x + 18)\) \(x + 6 = x + 6\) Since both of the sides of this equation equals to each other, we can say that this has infinite number of solutions.

OpenStudy (blackops2_143):

ok thx